OpenStudy (anonymous):

A motorist traveling at a constant velocity of 20 m/s passes a school-crossing corner where the speed limit is 10 m/s (about 22 mi/h). A police officer on a motorcycle stopped at the corner starts off in pursuit with a constant acceleration of 2.5 m/s^2. What is the officer's speed when she catches up with the car? What is the total distance of the officer has traveled at that point?

OpenStudy (anonymous):

set their two position functions equal to each other the function for the motorist is r=20t the function for the police car is r=(2.5/2)t^2 I assumed they started at the same time so 20t = (2.5/2)t^2 solve for t and then plug into the velocity function for police car and the position function of the police car

OpenStudy (anonymous):

Does this make sense to you?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

what position equation would you use to derive the function? my teacher is picky

OpenStudy (anonymous):

x=vo*t-(.5)*a*t^2